Status Report

NOAA GOES SATOPS Morning Report 27 September 2000

By SpaceRef Editor
September 27, 2000
Filed under

SATOPS MORNING REPORT – September 27, 2000

POES

No change in the status of the POES spacecraft.

The NOAA-16 Orbital Verification (OV) Test Team is continuing to report
routine operations.  The AMSU A1 and AMSU A2 instruments were turned
on and commanded to full scan mode yesterday without incident.  HRPT
receipt subsequent to the instrument turn-on activities was successful
and preliminary reports of AMSU output are positive.  The DCS system
was also turned on successfully using Data Recovery Unit 8.  Work
is proceeding to produce a consolidated interrogation schedule in anticipation
of NASA turning over health and safety operations of NOAA-16 on Friday,
September 29.  The SBUV instrument is scheduled to be turned on today. 
On-board computer command Macros to support OV communication testing are
expected to be loaded today as well.

The Fairbanks 13m “B” antenna was taken off-line at 1900Z on Tuesday. 
The downtime is expected to be 5 to 6 weeks and is necessary to perform
pre-winter modifications to the antenna.  POES and DMSP scheduling
is involved in this activity and will re-plan to take into this antenna
downtime into account.

POES operations were nominal over the past 24 hours.

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 24 hours.

GOES

No change in the status of the GOES spacecraft.

GOES-8 (East) Operations:

The GOES-8 fall eclipse season occurs daily between 0414-0526Z and continues
through October 15th.  GOES-8 images cancelled for Eclipse/KOZ – 0345,
0401, 0409, 0415, 0431, 0439, 0445, 0501, 0509, 0515, 0531, 0539, 0545Z
and 0615Z.  GOES-8 soundings cancelled for Eclipse/KOZ – 0346, 0420,
0446, 0520, 0546 and 0620Z.  Full disk coverage during GOES-10 KOZ:
0815Z – 1044Z.

At 1221Z on Tuesday, September 26, a GOES-8 frame break was observed
and attributed to loss of signal synchronization.

GOES-10 (West) Operations:

The GOES-10 fall eclipse season occurs daily between 0816-0930Z and
continues through October 16th.  GOES-10 images cancelled for KOZ/Eclipse:
0800, 0815, 0822, 0830, 0845, 0851, 0900, 0930, 1021 and 1030Z.  GOES-10
soundings cancelled for KOZ/Eclipse: 0801, 0824, 0901, 0924 and 1024Z. 
Full disk coverage during GOES-8 eclipse 0430Z – 0629Z.

At 1130Z on Tuesday, September 26, a GOES-10 frame break was observed
and attributed to high BER (unknown source).

At 1546Z on Tuesday, the HR2 antenna was configured for prime support
due to HR1 modifications.

GOES-10 WEFAX had an outage around 1547Z to 1610Z that was attributed
to a bad patch panel at the CDA (discovered during the HR1 — HR2 antenna
swap).  An additional GOES-10 WEFAX outage was reported between 1647Z
to 1820Z.  The reason is currently unknown, but it was noted that
the event coincided with GVAR troubleshooting between IPD and WWB (NSC).

At 1734Z on Tuesday, September 26, a GOES-10 frame break was observed
and attributed to loss of signal synchronization.

The 2245Z PACUS was dropped during Comm Controller failover activities.

Storage Mode Spacecraft Operations:

The GOES-9 fall eclipse season occurs daily between 0612-0724Z and continues
through October 15th.  GOES-9 remains in the Z-Axis Precession (ZAP)
storage mode.

The GOES-11 fall eclipse season occurs daily between 0619-0732Z and
continues through October 16th.  GOES-11 remains in the Z-Axis Precession
(ZAP) storage mode.

GOES operations were nominal over the past 24 hours.

 

Questions or comments to twalsh@nesdis.noaa.gov

SpaceRef staff editor.